# Everybody Prune Now: Structured Pruning of LLMs with only Forward Passes

### v2

@@ Appendix C Varying the pruning fraction per-iteration
Appendix D Varying the number of calibration data points for pruning
All Bonsai hyper-parameters are the same as Appendix A.2 except we vary $\mathrm{ns}_{\mathrm{data}}$ and $p_{\mathrm{iter}}=0.1$ to speed up pruning.
@@ Appendix D Varying the number of calibration data points for pruning
Appendix E Post-pruning adaptation
For this experiment, All Bonsai hyper-parameters are the same as Appendix A.2.
Appendix F Impact of prior
For this experiment, All Bonsai hyper-parameters are the same as Appendix A.2 except we vary $\rho$
F.1 $\rho$ is Activation Magnitude
@@
Appendix G How many perturbative samples are reasonable?
For this experiment, All Bonsai hyper-parameters are the same as Appendix A.2 except $p_{\mathrm{iter}}=0.1$ to speed up pruning.
@@ Appendix G How many perturbative samples are reasonable?
Appendix H Mistral-7B Experiment Details
In addition to the primary experiments on the LLaMA and Phi-2 models, supplementary experiments were performed on the Mistral-7B (Jiang et al., 2023) model in comparison with Wanda results on the stated model. We apply Bonsai with the same hardware and configuration settings as used for the LLaMA and Phi-2 experiments. We target different pruning fractions (0.05, 0.1, and 0.2) across different numbers of samples and masks per iteration to evaluate the method’s performance under varying sparsity conditions.
The Mistral-7B model architecture differs from the LLaMA architecture in its use of group query attention and sliding window attention in lieu of the standard self-attention used in most transformer-based models like LLaMA (Jiang et al., 2023). We factor these differences into consideration in the implementation of Bonsai for Mistral. For the experiments that produced the results below, all Bonsai hyper-parameters are the same as Appendix A.2.
@@ Appendix H Mistral-7B Experiment Details
We further investigate the pruning habits of Bonsai by examining the pruned layers of Mistral, as shown in Figure 4. We notice a recurring theme: when an attention layer is significantly altered, it leads to compensation in the next layers within the sequence. This adaptive behavior, termed the ”Hydra effect” by McGrath et al. (2023), implies that the layers within a language model interact in a way that changes in one layer prompt adjustments in another. McGrath et al. (2023) specifically mentioned that when one attention layer was removed from a language model, the model was still able to self-repair and produce similar outputs; but it did so by relying more heavily on other layers.
Figure 4: Mistral’s pruned attention layers. The heavily pruned layers are usually preceded by or sandwiched between lightly-pruned layers, exhibiting the self-repairing ”Hydra effect” (McGrath et al., 2023).

### v3

@@ Appendix E Varying the pruning fraction per-iteration
Appendix F Varying the number of calibration data points for pruning
All Bonsai hyper-parameters are the same as Appendix C.3 except we vary $\mathrm{ns}_{\mathrm{data}}$ and $p_{\mathrm{iter}}=0.1$ to speed up pruning.
@@ Appendix F Varying the number of calibration data points for pruning
Appendix G Post-pruning adaptation
For this experiment, All Bonsai hyper-parameters are the same as Appendix C.3.
Appendix H Impact of prior
For this experiment, All Bonsai hyper-parameters are the same as Appendix C.3 except we vary $\rho$.
H.1 $\rho$ is Activation Magnitude
@@
H.3 $\rho$ is Fluctuation-based (FLAP) An et al. (2024)
MLP / Fully Connected Module: Let $d$ be the intermediate dimension of the MLP to be pruned. Let $W\in\mathbb{R}^{d\times o}$ be the output projection matrix for the MLP. For any data-sample sequence $b$, we compute the sample variance of each input feature across batches and weight it with the squared norm of the corresponding column of the weight matrix:
$\left(\rho\in\mathbb{R}^{d}\right)\propto S_{:,j}=\frac{1}{N-1}\sum_{n=1}^{N}(X_{n,j,:}-\bar{X}_{:,j,:})^{2}\cdot||W_{:,j}||_{2}^{2}$ (7)
where $\bar{X}_{:,j,:}$ represents the average of the $j$-th channel for all samples, and $||W_{:,j}||_{2}^{2}$ denotes the squared norm of the $j$-th column of the weight matrix.
Self-Attention Module: We compute the fluctuation at the head level, weighted by the corresponding weights in the output projection matrix:
$\left(\rho\in\mathbb{R}^{h}\right)\propto S_{:,j}=\frac{1}{N-1}\sum_{n=1}^{N}(X_{n,j,:}-\bar{X}_{:,j,:})^{2}\cdot||W_{:,j}||_{2}^{2}$ (8)
For our improved fluctuation metric (fluct.2.0), we track the mean and second moment separately over multiple batches, computing variance as:
$\mathrm{Var}(X)=\mathbb{E}[X^{2}]-\mathbb{E}[X]^{2}$ (9)
This approach provides greater numerical stability when accumulating statistics across multiple forward passes.
Appendix I How many perturbative samples are reasonable?
For this experiment, All Bonsai hyper-parameters are the same as Appendix C.3 except $p_{\mathrm{iter}}=0.1$ to speed up pruning.
@@ Appendix I How many perturbative samples are reasonable?
Appendix J Mistral-7B Experiment Details
In addition to the primary experiments on the LLaMA and Phi-2 models, supplementary experiments were performed on the Mistral-7B Jiang et al. (2023) model in comparison with Wanda results on the stated model. We apply Bonsai with the same hardware and configuration settings as used for the LLaMA and Phi-2 experiments. We target different pruning fractions (0.05, 0.1, and 0.2) across different numbers of samples and masks per iteration to evaluate the method’s performance under varying sparsity conditions.
The Mistral-7B model architecture differs from the LLaMA architecture in its use of group query attention and sliding window attention in lieu of the standard self-attention used in most transformer-based models like LLaMA Jiang et al. (2023). We factor these differences into consideration in the implementation of Bonsai for Mistral. For the experiments that produced the results below, all Bonsai hyper-parameters are the same as Appendix C.3.
@@
We further investigate the pruning habits of Bonsai by examining the pruned layers of Mistral, as shown in Figure 6. We notice a recurring theme: when an attention layer is significantly altered, it leads to compensation in the next layers within the sequence. This adaptive behavior, termed the ”Hydra effect” by (McGrath et al., 2023), implies that the layers within a language model interact in a way that changes in one layer prompt adjustments in another. (McGrath et al., 2023) specifically mentioned that when one attention layer was removed from a language model, the model was still able to self-repair and produce similar outputs; but it did so by relying more heavily on other layers.
Figure 6: Mistral’s pruned attention layers. The heavily pruned layers are usually preceded by or sandwiched between lightly-pruned layers, exhibiting the self-repairing ”Hydra effect” McGrath et al. (2023).
Table 16: Varying $p_{\mathrm{iter}}$. Wikitext-2 perplexity of LLaMA-2 7B pruned to 50% sparsity. See Appendix E for experiment details.
$p_{\mathrm{iter}}=0.05$ $p_{\mathrm{iter}}=0.1$ $p_{\mathrm{iter}}=0.2$
w/o Adapt $19.47$ $61.63$ $209.44$
w Adapt $8.89$ $9.15$ $9.57$